# Lattice QCD Determination of the Collins–Soper Kernel in the Continuum and Physical Mass Limits

### v1

@@ Abstract
The Collins-Soper (CS) kernel governs the rapidity evolution of transverse-momentum-dependent (TMD) parton distributions, a cornerstone for QCD factorization and linking nucleon structure data across scales. Its nonperturbative behavior at large transverse separations ($b_{\perp}$) remains weakly constrained due to phenomenological model dependencies. We present a first-principles determination of the CS kernel at the continuum limit and physical pion mass from lattice QCD in the large-momentum effective theory framework. Using (2+1)-flavor configurations (lattice spacings a=(0.052a=(0.0520.1050.105) fm, and pion mass $m_{\pi}\approx(136,230,300,320)$ MeV), we simulating the nonlocal equal-time correlation function and extract the quasi-TMD wave functions. Taking into account systematic improvements including hypercubic smearing, nonperturbative renormalization, and a $b_{\perp}$-unexpanded matching kernel, we obtain the CS kernel at the continuum, chiral, and infinite-momentum limits. Our results are determined up to $b_{\perp}\sim 1$ fm, with controllable uncertainties, and agree with perturbative QCD at small $b_{\perp}$ and global TMD phenomenological extractions. We conduct a global analysis integrated with phenomenological fits and demonstrate the impact of our results on such fits. This work yields the most precise nonperturbative constraint on the CS kernel’s long-distance behavior from Lattice QCD, which not only bridges Lattice QCD, perturbation theory, and nucleon structure experiments for TMD studies, but also boosts the utility of our constraint for future global TMD analyses.
@@ I Introduction
Extensive experimental efforts over the past decades, including SIDIS at HERMES, COMPASS, and Jefferson Lab, as well as Drell–Yan and W/Z production at RHIC and the LHC, have provided rich information on TMD-sensitive observables HERMES:2009lmz; JeffersonLabHallA:2011ayy; COMPASS:2012dmt; HERMES:2012uyd; Anselmino:2013lza; STAR:2015vmv; COMPASS:2016led; Anselmino:2016uie; HERMES:2020ifk; Liu:2021boj; Cruz-Martinez:2023sdv; ATLAS:2023lsr; Isaacson:2023iui; COMPASS:2023vqt; Niemiec:2024eie. Global phenomenological analyses Scimemi:2019cmh; Bacchetta:2022awv; Moos:2023yfa; Bacchetta:2024qre; Kang:2024dja; Martinez:2024mou; Moos:2025sal; Bacchetta:2025ara have extracted the CS kernel from these data, revealing consistent results in the perturbative region but leaving sizable uncertainties in the nonperturbative domain due to model dependencies and parametrization assumptions.
@@
However, existing calculations of the CS kernel were limited in precision, particularly at large transverse separations, where the signal rapidly deteriorates, making it difficult to access the long-range correlations relevant to nucleon-size scales. A recent study Avkhadiev:2024mgd achieved the first multi-spacing calculation at physical pion mass, while the limited precision of the small lattice spacing data necessitated an extrapolation through parametrization. In the present work, we employ a set of ensembles with multiple lattice spacings and pion masses, together with substantially improved statistical precision. This allows us to perform controlled continuum and chiral extrapolations and to determine the CS kernel directly from first principles up to transverse separations of $b_{\perp}\simeq 1fm$. Our results thus provide the most precise and model-independent determination of the long-distance behavior of the CS kernel to date, offering a solid nonperturbative foundation for future global TMD analyses.
@@ III.2.1 Correlation Functions
Since the Dirac structure of pseudoscalar meson interpolators at the source does not influence the matrix elements, we choose $\Gamma_{2}=(\gamma^{t}\gamma_{5}+\gamma^{z}\gamma_{5})/2$ Ji:2024hit. At the sink, as mentioned earlier, we select $\Gamma_{1}=\gamma^{t}\gamma_{5}$, which will projects onto the leading-twist light-cone operator $\gamma^{+}\gamma_{5}$ in the large-momentum limit. Employing the reduction formula, the parametrized form of the two-point correlation function can be expressed as
@@ III.3 CS Kernel from Quasi-TMDWFs
$K(b_{\perp},\mu;x,P_{1}^{z},P_{2}^{z})$
$\qquad=K(b_{\perp},\mu)+A(x,b_{\perp},\mu)\left[\frac{1}{(P_{1}^{z})^{2}}-\frac{1}{(P_{2}^{z})^{2}}\right],$ (22)
which is motivated by the power corrections of the physical TMDWF. The term $A(x,b_{\perp},\mu)$ encodes higher-power effects that are enhanced near the endpoints of $x$. One can see from Fig. 14 that, the ratios from different momentum pairs are mutually consistent, indicating that the chosen momenta are sufficiently large to suppress higher-power contributions. Following Eq.(22), we carry out the extrapolation in the medial region $x\in[0.3,~0.7]$, where the power corrections are minimized, and obtain the leading-power CS kernel $K(b_{\perp},\mu)$. The resulting band is shown in gray in Fig. 14.
@@ III.4 Estimation of Systematic Uncertainties and Final Results
Figure 18: Left panel: comparison of our result $K(b_{\perp},\mu)$ the lattice calculations by LPC22 LatticePartonLPC:2022eev, ASWZ23 Avkhadiev:2023poz, DWF24 Bollweg:2024zet, PionCG25 Bollweg:2025iol, as well as the perturbative calculations up to three loops Li:2016ctv; Moch:2017uml. Right panel: comparison of our result with phenomenological extractions from SV19 Scimemi:2019cmh, ART23 Moos:2023yfa, IFY23 Isaacson:2023iui, EEC24 Kang:2024dja, ASWZ24 Avkhadiev:2024mgd, MAPNN25 Bacchetta:2025ara), ART25 Moos:2025sal.
@@
    LaMET requires sufficiently large boosted hadron momenta so that power corrections $\propto 1/\left(P^{z}\right)^{2}$ are under control. In practice, only finite momenta are accessible, and we extrapolate to $P^{z}\to\infty$ using Eq.(22) to remove leading $1/\left(P^{z}\right)^{2}$ effects. The associated systematic uncertainty is estimated by comparing the extrapolated value $K(b_{\perp},\mu)$ with the result obtained at the largest available lattice momentum and taking the absolute difference as $\delta^{P^{z}}_{\mathrm{sys}}$
@@
In addition, we also compare with results obtained from global fits, as shown in the right panel of Fig. 18. Specifically, we include SV19 Scimemi:2019cmh, ART23 Moos:2023yfa, IFY23 Isaacson:2023iui, EEC24 Kang:2024dja, MAPNN25 Bacchetta:2025ara) and ART25 Moos:2025sal. Such global analyses typically achieve smaller uncertainties than raw lattice determinations by imposing model parametrizations (see, e.g., Ref. Avkhadiev:2024mgd). Remarkably, without model parametrizations for the $b_{\perp}$ dependence, our determination attains comparable precision, particularly at large $b_{\perp}$. The agreement between our results and the global fits indicates full compatibility with phenomenological analyses within uncertainties.
@@ III.5 Some attempted global Fits using Phenomenology and Lattice Inputs
Only Phenomenological extractions $0.0621(31)$ $2.34$
Model 1 Lattice QCD $0.0642(42)$ $0.35$
Joint Analysis $0.0628(25)$ $1.49$
Only Phenomenological extractions $0.1613(40)$ $2.71$
Model 2 Lattice QCD $0.1451(65)$ $0.59$
Joint Analysis $0.1576(34)$ $1.89$
Only Phenomenological extractions 0.1702 (86) 2.77
Model 3 Lattice QCD 0.182 (11) 0.36
Joint Analysis 0.1745 (68) 1.75
@@
    Lattice only: fit using recent lattice calculations (DWF24 Bollweg:2024zet, PionCG25 Bollweg:2025iol, LPC25);
@@ IV Summary and outlook
Finally, through a combined extrapolation across all four ensembles, we achieved, for the first time, a determination of the CS kernel up to $1~\mathrm{fm}$ in the continuum limit and at the physical pion mass, without relying on model assumptions. All major sources of systematic uncertainty have been carefully analyzed and quantified, with their contributions explicitly presented. Our results are in good agreement with both previous lattice QCD calculations and phenomenological extractions from experimental data, demonstrating the reliability of this first-principles approach.
@@ Appendix A Self Renormalization
Figure 19: The result of joint fit for the self renormalization, taking $b=\{0.15,0.2,0.25,0.3\}~\mathrm{fm}$ and $z=0.1~\mathrm{fm}$ as an example(left panel). Matching result of the $g(z,b_{\perp})$ to the $\overline{\text{MS}}$ scheme with HYP=1(right panel).
@@ D.2 Results for self renormalization
More results for the self-renormalization on different HYP smearing steps are given in Fig. 22, 23 , 24 and 25.
Figure 22: The result of joint fit for the self renormalization with HYP=0, $b=\{0.15,0.2,0.25,0.3\}~\mathrm{fm}$ and $z=\{0.15,0.2,0.25\}~\mathrm{fm}$ .
Figure 23: The result of joint fit for the self renormalization with HYP=1, $b=\{0.15,0.2,0.25,0.3\}~\mathrm{fm}$ and $z=\{0.15,0.2,0.25\}~\mathrm{fm}$ .
Figure 24: The result of joint fit for the self renormalization with HYP=2, $b=\{0.15,0.2,0.25,0.3\}~\mathrm{fm}$ and $z=\{0.15,0.2,0.25\}~\mathrm{fm}$ .
@@ D.3 Results for fits to the bare quasi-TMDWF matrix elements
More results for the fits to the bare quasi-TMDWF matrix elements on different ensembles are given in Fig. 26,  27,  28 and  29.
@@ D.4 Results for renormalized quasi-TMDWF matrix elements
More results for the renormalized quasi-TMDWF matrix elements on different ensembles are given in Fig. 30,  31,  32 and  33.
@@ D.5 Results for quasi-TMDWFs
More results for the quasi-TMDWFs on different ensembles are given in Fig. 34,  35,  36 and  37.
@@ D.6 Results for CS kernel
More results for the CS kernel on different ensembles are given in Fig. 38,  39,  40 and  41.

### v2

@@ Abstract
The Collins-Soper (CS) kernel governs the rapidity evolution of transverse-momentum-dependent (TMD) parton distributions, a cornerstone for QCD factorization and linking nucleon structure data across scales. Its nonperturbative behavior at large transverse separations ($b_{\perp}$) remains weakly constrained due to phenomenological model dependencies. We present a first-principles determination of the CS kernel at the continuum limit and physical pion mass from lattice QCD in the large-momentum effective theory framework. Using (2+1)-flavor configurations (lattice spacings $a\in[0.052,0.105]$ fm, and pion mass $m_{\pi}\approx(136,230,300,320)$ MeV), we simulating the nonlocal equal-time correlation function and extract the quasi-TMD wave functions. Taking into account systematic improvements including hypercubic smearing, nonperturbative renormalization, and a $b_{\perp}$-unexpanded matching kernel, we obtain the CS kernel at the continuum, chiral, and infinite-momentum limits. Our results are determined up to $b_{\perp}\sim 1$ fm, with controllable uncertainties, and agree with perturbative QCD at small $b_{\perp}$ and global TMD phenomenological extractions. We conduct a global analysis integrated with phenomenological fits and demonstrate the impact of our results on such fits. This work yields the most precise nonperturbative constraint on the CS kernel’s long-distance behavior from Lattice QCD, which not only bridges Lattice QCD, perturbation theory, and nucleon structure experiments for TMD studies, but also boosts the utility of our constraint for future global TMD analyses.
@@ I Introduction
Extensive experimental efforts over the past decades, including SIDIS at HERMES, COMPASS, and Jefferson Lab, as well as Drell–Yan and W/Z production at RHIC and the LHC, have provided rich information on TMD-sensitive observables HERMES:2009lmz; JeffersonLabHallA:2011ayy; COMPASS:2012dmt; HERMES:2012uyd; Anselmino:2013lza; STAR:2015vmv; COMPASS:2016led; Anselmino:2016uie; HERMES:2020ifk; Liu:2021boj; Cruz-Martinez:2023sdv; ATLAS:2023lsr; Isaacson:2023iui; COMPASS:2023vqt; Niemiec:2024eie. Global phenomenological analyses Scimemi:2019cmh; Bacchetta:2022awv; Moos:2023yfa; Bacchetta:2024qre; Kang:2024dja; Martinez:2024mou; Moos:2025sal; Bacchetta:2025ara; Camarda:2025lbt have extracted the CS kernel from these data, revealing consistent results in the perturbative region but leaving sizable uncertainties in the nonperturbative domain due to model dependencies and parametrization assumptions.
@@
However, existing calculations of the CS kernel were limited in precision, particularly at large transverse separations, where the signal rapidly deteriorates, making it difficult to access the long-range correlations relevant to nucleon-size scales. A recent study Avkhadiev:2024mgd performed the first continuum-extrapolated calculation at the physical pion mass using multiple lattice spacings in the range $a\in[0.09,0.15]~\mathrm{fm}$. Their continuum extrapolation, however, relied on an explicit parametrization of the $b_{\perp}$ dependence of the CS kernel, because discretization effects and power corrections could not be disentangled simultaneously at each $b_{\perp}$ within the available data constraints. In this work, we employ ensembles with multiple lattice spacings spanning $a\in[0.052,0.105]~\mathrm{fm}$ and multiple pion masses, including the physical one, together with substantially improved statistical precision. This enables controlled continuum and chiral extrapolations and allows us to determine the CS kernel from first principles up to transverse separations of $b_{\perp}\simeq 1~\mathrm{fm}$. Our results therefore provide, to date, the most precise determination of the long-distance CS kernel and the first continuum-extrapolated extraction achieved without introducing the explicit parametrization of its $b_{\perp}$ dependence, thereby offering a robust nonperturbative input for future global TMD analyses.
@@ III.2.1 Correlation Functions
Since the Dirac structure of pseudoscalar meson interpolators at the source does not influence the matrix elements, we choose $\Gamma_{2}=(\gamma^{t}\gamma_{5}+\gamma^{z}\gamma_{5})/2$ Zhang:2025hyo. At the sink, as mentioned earlier, we select $\Gamma_{1}=\gamma^{t}\gamma_{5}$, which will projects onto the leading-twist light-cone operator $\gamma^{+}\gamma_{5}$ in the large-momentum limit. Employing the reduction formula, the parametrized form of the two-point correlation function can be expressed as
@@ III.3 CS Kernel from Quasi-TMDWFs
$K(b_{\perp},\mu;x,P_{1}^{z},P_{2}^{z},a,m_{\pi})=K(b_{\perp},\mu;a,m_{\pi})$
$\qquad+A(b_{\perp},\mu;x,a,m_{\pi})\left[\frac{1}{(P_{1}^{z})^{2}}-\frac{1}{(P_{2}^{z})^{2}}\right],$ (22)
which is motivated by factorization analysis of the TMDWF in Eq. (3). The coefficient $A(b_{\perp},\mu;x,a,m_{\pi})$ parametrizes next-to-leading power effects, which can be numerically enhanced at the endpoint regions in $x$.
The extractions of $K(b_{\perp},\mu;a,m_{\pi})$ in the infinite-momentum limit are shown in Fig. 14. We find that the results obtained from different momentum pairs are mutually consistent, indicating that the chosen hadron momenta are sufficiently large and that residual higher-power contributions are well suppressed. Based on Eq. (22), we therefore perform the infinite-momentum extrapolation in the intermediate-$x$ window, $x\in[0.3,0.7]$, where power corrections are expected to be minimized. The resulting leading power CS kernel $K(b_{\perp},\mu;a,m_{\pi})$ are presented as the gray band in Fig. 14.
@@
Beyond the leading $1/(P^{z})^{2}$ corrections included in Eq. (22), there also may exist potential lattice-spacing–dependent discretization corrections proportional to $aP^{z}$ Avkhadiev:2024mgd. Such contributions may originate from discretization artifacts at large hadron momentum and could, in principle, affect the infinite-momentum extrapolation.
Beyond the leading $1/(P^{z})^{2}$ corrections included in Eq. (22), additional lattice-spacing–dependent power corrections proportional to $aP^{z}$ may also arise, as suggested in Ref. Avkhadiev:2024mgd. Such terms can originate from discretization artifacts at large hadron momenta and could, in principle, affect the infinite-momentum extrapolation in Eq. (22). In the present analysis, we observe that these corrections are small for two reasons. First, as shown in Fig. 2, the boosted hadron energies on all ensembles are consistent with the continuum dispersion relation over the range of momenta used, indicating that discretization effects that grow with $P^{z}$ are well under control. Secondly, as illustrated in the left panel of Fig. 15, the CS kernels extracted at the same momentum combinations but at different lattice spacings $a$ agree within uncertainties, providing direct evidence that the residual $aP^{z}$-dependent effects are swamped by the statistical errors.
To further validate this assumption, we combine the extrapolation ansätze in Eqs. (22) and (24), and explicitly include a mixed term proportional to $aP^{z}$ to parameterize lattice-spacing–dependent power corrections suggested in Ref. Avkhadiev:2024mgd,
$K(b_{\perp},\mu;x,P^{z}_{1},P^{z}_{2},a,m_{\pi})=K(b_{\perp},\mu)$
$\qquad+A(x,b_{\perp},\mu)\!\left[\frac{1}{(P^{z}_{1})^{2}}-\frac{1}{(P^{z}_{2})^{2}}\right]$
$\qquad+a^{2}B(b_{\perp},\mu)+(m_{\pi}^{2}-m_{\pi,\mathrm{phy}}^{2})C(b_{\perp},\mu)$
$\qquad+D(b_{\perp},\mu)\,a(P^{z}_{1}-P^{z}_{2}),$ (25)
where the term proportional to $D(b_{\perp},\mu)$ is introduced to capture potential discretization artifacts that grow linearly with the hadron momentum.
The comparison of the extrapolated result of $K(b_{\perp},\mu)$ from the fitting strategies i) a chained fit based on Eqs. (22) and (24) and ii) a joint fit from Eq. (25) are displayed in Fig. 17. One can see that within the current statistical precision, the results for two strategies are consistent within uncertainties. In particular, no statistically significant $aP^{z}$ term can be resolved, indicating that its contribution cannot be reliably disentangled from the leading power corrections with the present data. For the remainder of this work, we adopt the results obtained from the chained fitting strategy for our final analysis.
Figure 17: Comparison of the extrapolated CS kernel $K(b_{\perp},\mu)$ obtained with two fitting strategies: (i) a chained fit based on Eqs. (22) and (24) and (ii) a joint fit using Eq. (25). Only statistical uncertainties are shown.
@@ III.4 Estimation of Systematic Uncertainties and Final Results
Figure 19: Left panel: comparison of our result $K(b_{\perp},\mu)$ the lattice calculations by LPC22 LatticePartonLPC:2022eev, ASWZ23 Avkhadiev:2023poz, DWF24 Bollweg:2024zet, PionCG25 Bollweg:2025iol, as well as the perturbative calculations up to three loops Li:2016ctv; Moch:2017uml. Right panel: comparison of our result with phenomenological extractions from SV19 Scimemi:2019cmh, ART23 Moos:2023yfa, IFY23 Isaacson:2023iui, EEC24 Kang:2024dja, ASWZ24 Avkhadiev:2024mgd, MAPNN25 Bacchetta:2025ara, ART25 Moos:2025sal, CFR25 Camarda:2025lbt.
@@
    LaMET requires sufficiently large hadron boosts such that power corrections of order $1/(P^{z})^{2}$ are under control. In practice, only finite momenta are accessible, and we therefore extrapolate to $P^{z}\to\infty$ using Eq. (22) to remove the leading $1/(P^{z})^{2}$ effects. We estimate the associated systematic uncertainty by comparing the extrapolated value $K(b_{\perp},\mu)$ with the result obtained at the largest available lattice momentum, and take the absolute difference as $\delta^{P^{z}}_{\mathrm{sys}}$.
  5. 5.
    $\delta^{aP^{z}}_{\mathrm{sys}}$: from lattice-spacing-dependent power corrections.
    As discussed in Sec. III.3, results from the joint-fit strategy based on Eq. (25) are consistent with those from the chained-fit strategy. For a conservative error budget, we adopt the latter one in the final analysis, which allows us to estimate the systematic uncertainties from the continuum extrapolation and the infinite-momentum extrapolation separately. Accordingly, we do not introduce an independent uncertainty $\delta^{aP^{z}}_{\mathrm{sys}}$; instead, any residual $aP^{z}$-dependent effects are treated as subleading and are effectively covered by the systematic variations already included in $\delta^{P^{z}}_{\mathrm{sys}}$ and $\delta^{a}_{\mathrm{sys}}$.
@@
In addition, we compare our results with those from global analyses, as shown in the right panel of Fig. 19. Specifically, we include SV19 Scimemi:2019cmh, ART23 Moos:2023yfa, IFY23 Isaacson:2023iui, EEC24 Kang:2024dja, MAPNN25 Bacchetta:2025ara, ART25 Moos:2025sal, and CFR25 Camarda:2025lbt. Usually, global analyses typically achieve smaller uncertainties than direct lattice determinations by imposing explicit model parametrizations, in particular for the $b_{\perp}$ dependence (see, e.g., Ref. Avkhadiev:2024mgd). Remarkably, our determination reaches a comparable precision, especially at large $b_{\perp}$, without introducing a dedicated model parametrization for the $b_{\perp}$ dependence in the continuum-extrapolated CS kernel. In our analysis, the independence of a $b_{\perp}$ ansatz is enabled by the infinite-momentum extrapolation in Eq. (22), where we find no statistically significant sensitivity to the mixed discretization term proportional to $aP^{z}$ and therefore neglect it at the current level of precision. The resulting agreement between our CS kernel and the global-fit determinations demonstrates full compatibility with phenomenological extractions within uncertainties.
@@ III.5 Some attempted global Fits using Phenomenology and Lattice Inputs
Only Phenomenological extractions $0.0702(28)$ $2.12$
Model 1 Lattice QCD $0.0711(36)$ $0.80$
Joint Analysis $0.0705(22)$ $1.46$
Only Phenomenological extractions $0.1570(33)$ $3.34$
Model 2 Lattice QCD $0.1435(46)$ $1.02$
Joint Analysis $0.1521(27)$ $2.26$
Only Phenomenological extractions 0.1972 (81) 2.73
Model 3 Lattice QCD 0.210 (10) 0.97
Joint Analysis 0.1996 (64) 1.86
@@
    Lattice only: fit using recent lattice calculations (ASWZ23 Avkhadiev:2023poz, DWF24 Bollweg:2024zet, PionCG25 Bollweg:2025iol, LPC25);
@@ IV Summary and outlook
Finally, through a combined extrapolation across all four ensembles, we achieved, for the first time, a determination of the CS kernel up to $1~\mathrm{fm}$ in the continuum limit and at the physical pion mass without introducing any explicit parametrization of its $b_{\perp}$ dependence. All major sources of systematic uncertainty have been carefully analyzed and quantified, with their contributions explicitly presented. Our results are in good agreement with both previous lattice QCD calculations and phenomenological extractions from experimental data, demonstrating the reliability of this first-principles approach.
@@ Appendix A Self Renormalization
Figure 20: The result of joint fit for the self renormalization, taking $b_{\perp}=\{0.15,0.2,0.25,0.3\}~\mathrm{fm}$ and $z=0.1~\mathrm{fm}$ as an example(left panel). Matching result of the $g(z,b_{\perp})$ to the $\overline{\text{MS}}$ scheme with HYP=1(right panel).
@@ D.2 Results for self renormalization
More results for the self-renormalization on different HYP smearing steps are given in Fig. 23, 24 , 25 and 26.
Figure 23: The result of joint fit for the self renormalization with HYP=0, $b_{\perp}=\{0.15,0.2,0.25,0.3\}~\mathrm{fm}$ and $z=\{0.15,0.2,0.25\}~\mathrm{fm}$ .
Figure 24: The result of joint fit for the self renormalization with HYP=1, $b_{\perp}=\{0.15,0.2,0.25,0.3\}~\mathrm{fm}$ and $z=\{0.15,0.2,0.25\}~\mathrm{fm}$ .
Figure 25: The result of joint fit for the self renormalization with HYP=2, $b_{\perp}=\{0.15,0.2,0.25,0.3\}~\mathrm{fm}$ and $z=\{0.15,0.2,0.25\}~\mathrm{fm}$ .
@@ D.3 Results for fits to the bare quasi-TMDWF matrix elements
More results for the fits to the bare quasi-TMDWF matrix elements on different ensembles are given in Fig. 27,  28,  29 and  30.
@@ D.4 Results for renormalized quasi-TMDWF matrix elements
More results for the renormalized quasi-TMDWF matrix elements on different ensembles are given in Fig. 31,  32,  33 and  34.
@@ D.5 Results for quasi-TMDWFs
More results for the quasi-TMDWFs on different ensembles are given in Fig. 35,  36,  37 and  38.
@@ D.6 Results for CS kernel
More results for the CS kernel on different ensembles are given in Fig. 39,  40,  41 and  42.